# Weak expansion properties and a large deviation principle for coarse expanding conformal systems

### v2

@@ 5.2. Technical lemmas
$\inf\limits_{U\in\mathcal{U}_{n}}\operatorname{diam}U\geq C^{\prime\prime}\lambda^{n}.$
@@
$\operatorname{diam}U^{\prime}>\lambda\operatorname{diam}U\geq\lambda\inf\limits_{V\in\mathcal{U}_{n}}\operatorname{diam}V.$
@@
$\inf\limits_{V^{\prime}\in\mathcal{U}_{n+1}}\operatorname{diam}V^{\prime}\geq\lambda\inf\limits_{V\in\mathcal{U}_{n}}\operatorname{diam}V.$
@@
$\inf\limits_{V\in\mathcal{U}_{n}}\operatorname{diam}V\geq\lambda^{n}\inf\limits_{V^{0}\in\mathcal{U}_{0}}\operatorname{diam}\bigl{(}V^{0}\bigr{)}.$
@@
It should be noted that since $c_{n}\leq d_{n}$, we have $10<\lambda\leq\theta<$ 1, where the constant $\theta\in(0,1)$ is from Proposition 3.11.
@@
(5.6) $\operatorname{B}_{m}^{k}(U^{m})\coloneqq$ $\bigcup\mathcal{B}_{m}^{k}(U^{m})$
$=$ $U^{m}\cup\bigcup\bigl{\{}U^{m+k}\in\mathcal{U}_{m+k}:U^{m+k}\cap U^{m}\cap X\neq\emptyset\bigr{\}}.$
@@
    If $a,b\in U$, then $\lvert a-b\rvert\leq\operatorname{diam}U$.
@@
    ${\operatorname{diam}U^{\prime}}/{\operatorname{diam}U}\leq C^{\prime}\theta^{k},$
    and consequently $\lvert a-b\rvert\leq\bigl{(}C^{\prime}\theta^{k}+1\bigr{)}\operatorname{diam}U$.
@@
    If there exist $U^{\prime},\,U^{\prime\prime}\in\mathcal{U}_{m+k}$ with $U\cap U^{\prime}\cap X\neq\emptyset$ and $U\cap U^{\prime\prime}\cap X\neq\emptyset$, such that $a\in U^{\prime}$, $b\in U^{\prime\prime}$, then by a similar argument as in (ii), we get $\lvert a-b\rvert\leq 2\bigl{(}C^{\prime}\theta^{k}+1\bigr{)}\operatorname{diam}U$.
Therefore, $\operatorname{diam}\bigl{(}\operatorname{B}_{m}^{k}(U)\bigr{)}\leq 2\bigl{(}C^{\prime}\theta^{k}+1\bigr{)}\operatorname{diam}U<2(C^{\prime}+1)^{2}d_{0}\theta^{m}$. ∎
Assume that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4. We consider a sequence of open covers $\{\mathcal{W}_{n}\}_{n\in\mathbb{N}_{0}}$ of $X$, defined as
(5.8) $\mathcal{W}_{n}\coloneqq\{U\cap X:U\in\mathcal{U}_{n}\},$
and show some properties of subsets of the form
@@
Assume that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4. Denote $f_{X}\coloneqq f|_{X}$. Then for each $n\in\mathbb{N}_{0}$ and each $W^{n}\in\mathcal{W}_{n}$, if $W^{n}=U^{n}\cap X$ for some $U^{n}\in\mathcal{U}_{n}$, then
(5.10) $f_{X}^{-1}(W^{n})=f^{-1}(U^{n})\cap X=X\cap\bigcup\bigl{\{}U^{n+1}\in\mathcal{U}_{n+1}:f\bigl{(}U^{n}\bigr{)}=U^{n+1}\bigr{\}}.$
The lemma below shows that the subsets of the form
$A=\bigcap_{i=0}^{n}f_{X}^{-i}(W_{i}^{m})=\left\{x\in W_{0}^{m}:f_{X}^{i}(x)\in W^{m}_{i},\ i\in\{1,2,\dots,n\}\right\}$
can be covered by certain elements in $\mathcal{W}_{m+n}$.
First, for all $m\in\mathbb{N}_{0}$, $n\in\mathbb{N}$, and for an arbitrary choice of $U^{m}_{i}\in\mathcal{U}_{m}$ for each $i\in\{0,1,\dots,n\}$, we denote
@@
[…]
Then the following lemma holds.
@@
Assume that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4. Denote $f_{X}\coloneqq f|_{X}$. Then for all $m\in\mathbb{N}_{0}$, $n\in\mathbb{N}$, and an arbitrary choice of $U^{m}_{i}\in\mathcal{U}_{m}$ for each $i\in\{0,1,\dots,n\}$, we have
@@
[…]
@@
Assume that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4. Denote $\mathcal{W}_{n}\coloneqq\{U\cap X:U\in\mathcal{U}_{n}\}$ for each $n\in\mathbb{N}_{0}$. Then there exists a constant $T_{0}\geq 1$ such that for all $n\in\mathbb{N}_{0}$, $k\in\mathbb{N}$, and $W^{n}\in\mathcal{W}_{n}$, there exists $\mathcal{I}\subseteq\mathcal{W}_{n+k}$ with $\operatorname{card}(\mathcal{I})\leq(pT_{0})^{k}$ such that
@@
Fix arbitrary integer $n\geq N$ and $U^{n}\in\mathcal{U}_{n}$, and consider $\operatorname{B}_{n}^{1}(U^{n})$. We have
@@
for some $U^{0}\in\mathcal{U}_{0}$. Since $U^{n}$ is a preimage under $f^{n-N}$ of $f^{n-N}(U^{n})$, there exists a preimage under $f^{n-N}$ of $U^{0}$, denoted by $U^{n-N}\in\mathcal{U}_{n-N}$, such that $U^{n}\subseteq U^{n-N}$. For each $U^{n+1}\in\mathcal{B}_{n}^{1}(U^{n})\cap\mathcal{U}_{n+1}$, there exists $U^{\prime}\in\mathcal{U}_{n-N}$ such that $U^{n+1}\subseteq U^{\prime}$. Since $U^{n+1}\cap U^{n}\neq\emptyset$ and different preimages of $U^{0}$ are disjoint, we have $U^{\prime}=U^{n-N}$, thus $\operatorname{B}_{n}^{1}(U^{n})\subseteq U^{n-N}$.
Consider $f|_{U^{n-N}}\colon U^{n-N}\rightarrow U^{0}$. By [Deg], this map is at most $p$ to $1$, thus we have
$\operatorname{card}\left(\mathcal{B}_{n}^{1}(U^{n})\right)\leq p\cdot\operatorname{card}\bigl{(}\mathcal{B}_{N}^{1}(f^{n-N}(U^{n})\bigr{)}\bigr{)}.$
Let $T^{\prime}_{0}\coloneqq\max\limits_{0\leq i\leq N}\sup\bigl{\{}\operatorname{card}\bigl{(}\mathcal{B}_{i}^{1}(U)\bigr{)}:U\in\mathcal{U}_{i}\bigr{\}}$, then we have
$\operatorname{card}\left(\mathcal{B}_{n}^{1}(U^{n})\right)\leq pT^{\prime}_{0}.$
@@
Thus each element of $\mathcal{W}_{n}$ can be covered by at most $pT^{\prime}_{0}$ many elements in $\mathcal{W}_{n+1}$. Inductively, for every $k\in\mathbb{N}$, each element in $\mathcal{W}_{n}$ can be covered by $(pT^{\prime}_{0})^{k}$ many elements in $\mathcal{W}_{n+k}$. We have established the lemma under the additional assumption that $n\geq N$.
Finally, it is now easy to check that the lemma holds for general $n\in\mathbb{N}_{0}$ by setting $T_{0}\coloneqq 1+d^{N+1}$ since $T_{0}\geq\max\limits_{0\leq i\leq N+1}\operatorname{card}\mathcal{U}_{i}\geq T^{\prime}_{0}$. ∎
5.3. Proof of Theorems A and B
In this subsection, we will establish the asymptotic $h$-expansiveness of the map $f|_{X}\colon X\rightarrow X$. Our strategy is to prove the result under slightly stronger assumptions, then we show that for some $N\in\mathbb{N}$, $(f|_{X})^{N}$ satisfies the stronger assumptions, resulting in $h^{*}\left((f|_{X})^{N}\right)=Nh^{*}(f|_{X})=0$.
First, we note that $X$ is indeed a compact metric space. By $X=\bigcap_{n}\overline{\mathfrak{X}_{n}}$, $X$ is closed in $\overline{\mathfrak{X}_{1}}$. Since $\overline{\mathfrak{X}_{1}}$ is compact, $X$ is compact in $\overline{\mathfrak{X}_{1}}$, thus compact with respect to its own subspace topology.
@@ 5.3. Proof of Theorems A and B
Here $\mathcal{W}_{n}\coloneqq\{U\cap X:U\in\mathcal{U}_{n}\}$ for each $n\in\mathbb{N}_{0}$ is defined (5.8).
@@
Let $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4. Denote $f_{X}\coloneqq f|_{X}$. Assume that for each $n\in\mathbb{N}_{0}$, $\mathcal{U}_{n+1}$ is a refinement of $\mathcal{U}_{n}$. Then $f|_{X}\colon X\rightarrow X$ is asymptotically h-expansive.
@@
By Lemma 5.11, $\{\mathcal{W}_{m}\}_{m\in\mathbb{N}_{0}}$ (defined in (5.8)) is a refining sequence of open cover of $X$.
@@
[…]
$\mathcal{L}_{k}\coloneqq\left\{f^{k}(U):U\in E_{m}(U^{m}_{0},\dots,U^{m}_{n-1};U^{m}_{n})\right\}.$
@@
We can check that $\{M_{m}\}_{m\in\mathbb{N}_{0}}$ is increasing to +$\infty$+ ∞ and that
@@
$\operatorname{diam}\operatorname{B}_{m}^{n-k}(U^{m}_{k})<C^{\prime\prime}\lambda^{M_{n}}/(2K)\leq\delta_{M_{m}},$
where $\delta_{M_{m}}$ is the Lebesgue number of $\mathcal{U}_{M_{m}}$, which is from Proposition 3.10 (iv). Thus there exists $\widehat{U}_{k}\in\mathcal{U}_{M_{m}}$ such that
@@
For each $0\leq k\leq n$ and each $0\leq t\leq\min\{n-k,M_{m}\}$, consider $f^{t}|_{\widehat{U}_{k}}\colon\widehat{U}_{k}\rightarrow f^{t}(\widehat{U}_{k})$. By [Deg], the map is at most $p$ to 1, so $\operatorname{card}(\mathcal{L}_{k})\leq p\cdot\operatorname{card}(\mathcal{L}_{k+t})$.
Inductively we have $\operatorname{card}(\mathcal{L}_{0})\leq p^{\frac{n}{M_{m}}+1}\cdot\operatorname{card}(\mathcal{L}_{n})$, which is equivalent to
(5.16) $\operatorname{card}\left(E_{m}(U^{m}_{0},\dots,U^{m}_{n-1};U^{m}_{n})\right)\leq p^{\frac{n}{M_{m}}+1}.$
Define $\mathcal{I}_{A}\coloneqq\{U\cap X:U\in E_{m}(U^{m}_{0},\dots,U^{m}_{n-1};U^{m}_{n})\}\subseteq\mathcal{W}_{m+n}$. Then $\operatorname{card}(\mathcal{I}_{A})\leq p^{\frac{n}{M_{m}}+1}$, and $A\subseteq\bigcup\mathcal{I}_{A}$.
@@
By Lemma 5.10, for each $W^{m+n}\in\mathcal{I}_{A}$ there exists $\mathcal{I}_{W^{m+n}}\subseteq\mathcal{W}_{l+n}$ with $\operatorname{card}(\mathcal{I}_{W^{m+n}})\leq(pT_{0})^{l-m}$ such that $W^{m+n}\subseteq\bigcup\mathcal{I}_{W^{m+n}}$, where $p$ is the constant in [Deg] and $T_{0}$ is from Lemma 5.10.
@@
For each $W^{l+n}\in\mathcal{W}_{l+n}$, suppose that $W^{l+n}=U^{l+n}\cap X$, $U^{l+n}\in\mathcal{U}_{l+n}$. By the hypothesis in the statement of Theorem 5.12, that $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ is refining, for each $j\in\{0,1,\dots,n\}$, there exists $U^{l+j}\in\mathcal{U}_{l+j}$ such that $U^{l+n}\subseteq U^{l+j}$. Since $U^{l+j}\subseteq f^{-j}\bigl{(}U^{l}_{j}\bigr{)}$ for some $U^{l}_{j}\in\mathcal{U}_{l}$, we have $U^{l+n}\subseteq U^{l+j}\subseteq f^{-j}\bigl{(}U^{l}_{j}\bigr{)}$. As a consequence
@@
$=$ $\bigcap_{j=0}^{n}(X\cap f^{-j}\bigl{(}U^{l}_{j})\bigr{)}=\bigcap_{j=0}^{n}f^{-j}\bigl{(}X\cap U^{l}_{j}\bigr{)}\in\bigvee_{j=0}^{n}f^{-j}(\mathcal{W}_{l}).$
By Steps 1 through 3, there exists $\mathcal{J}_{A}\subseteq\bigvee_{j=0}^{n}f_{X}^{-j}(\mathcal{W}_{l})$ with $\operatorname{card}(\mathcal{J}_{A})\leq p^{\frac{n}{M_{m}}+1}(pT_{0})^{l-m}$, such that $A\subseteq\mathcal{J}_{A}$. It follows that
$\min\biggl{\{}\operatorname{card}\xi_{A}:\xi_{A}\subseteq\bigvee_{j=0}^{n}f_{X}^{-j}(\mathcal{W}_{l}),\ A\subseteq\bigcup\xi_{A}\biggr{\}}\leq p^{\frac{n}{M_{m}}+1}(pT_{0})^{l-m}.$
@@
$\qquad\leq\log\bigl{(}p^{\frac{n}{M_{m}}+1}(pT_{0})^{l-m}\bigr{)}.$
@@
$\leq\lim_{m\to+\infty}\lim_{l\to+\infty}\lim_{n\to+\infty}\frac{1}{n}\log\left(p^{\frac{n-1}{M_{m}}+1}(pT_{0})^{l-m}\right)$
$=\lim_{m\to+\infty}\frac{\log p}{M_{m}}$
$=0.$

### v3

@@ 5.2. Technical lemmas
$\inf\limits_{U\in\mathcal{U}_{n}}\operatorname{diam}U\geqslant C^{\prime\prime}\lambda^{n}.$
@@
$\operatorname{diam}U^{\prime}>\lambda\operatorname{diam}U\geqslant\lambda\inf\limits_{V\in\mathcal{U}_{n}}\operatorname{diam}V.$
@@
$\inf\limits_{V^{\prime}\in\mathcal{U}_{n+1}}\operatorname{diam}V^{\prime}\geqslant\lambda\inf\limits_{V\in\mathcal{U}_{n}}\operatorname{diam}V.$
@@
$\inf\limits_{V\in\mathcal{U}_{n}}\operatorname{diam}V\geqslant\lambda^{n}\inf\limits_{V^{0}\in\mathcal{U}_{0}}\operatorname{diam}\bigl{(}V^{0}\bigr{)}.$
@@
It should be noted that since $c_{n}\leqslant d_{n}$, we have $0<\lambda\leqslant\theta<1$, where the constant $\theta\in(0,1)$ is from Proposition 3.11.
@@
(5.7) $\operatorname{B}_{m}^{k}(U^{m})\coloneqq\bigcup\mathcal{B}_{m}^{k}(U^{m})=U^{m}\cup\bigcup\bigl{\{}U^{m+k}\in\mathcal{U}_{m+k}:U^{m+k}\cap U^{m}\cap X\neq\emptyset\bigr{\}}.$
@@
    If $a,b\in U$, then $\lvert a-b\rvert\leqslant\operatorname{diam}U$.
@@
    ${\operatorname{diam}U^{\prime}}/{\operatorname{diam}U}\leqslant C^{\prime}\theta^{k},$
    and consequently $\lvert a-b\rvert\leqslant\bigl{(}C^{\prime}\theta^{k}+1\bigr{)}\operatorname{diam}U$.
@@
    If there exist $U^{\prime},\,U^{\prime\prime}\in\mathcal{U}_{m+k}$ with $U\cap U^{\prime}\cap X\neq\emptyset$ and $U\cap U^{\prime\prime}\cap X\neq\emptyset$, such that $a\in U^{\prime}$, $b\in U^{\prime\prime}$, then by a similar argument as in (ii), we get $\lvert a-b\rvert\leqslant 2\bigl{(}C^{\prime}\theta^{k}+1\bigr{)}\operatorname{diam}U$.
Therefore, $\operatorname{diam}\bigl{(}\operatorname{B}_{m}^{k}(U)\bigr{)}\leqslant 2\bigl{(}C^{\prime}\theta^{k}+1\bigr{)}\operatorname{diam}U<2(C^{\prime}+1)^{2}d_{0}\theta^{m}$. ∎
In what follows, we show some properties of subsets of the form
@@
Assume that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4. Then for each $n\in\mathbb{N}_{0}$ and each $W^{n}\in\mathcal{W}_{n}$, if $W^{n}=U^{n}\cap X$ for some $U^{n}\in\mathcal{U}_{n}$, then
(5.10) $f_{X}^{-1}(W^{n})=f^{-1}(U^{n})\cap X=X\cap\bigcup\bigl{\{}U^{n+1}\in\mathcal{U}_{n+1}:f\bigl{(}U^{n+1}\bigr{)}=U^{n}\bigr{\}}.$
We construct a class of open covers for subsets of the form $A=\bigcap_{i=0}^{n}f_{X}^{-i}(W_{i}^{m})$. For all $m\in\mathbb{N}_{0}$, $n\in\mathbb{N}$, and for an arbitrary choice of $U^{m}_{i}\in\mathcal{U}_{m}$ for each $i\in\{0,1,\dots,n\}$, we denote
@@
 $\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ for each }i\in\{0,\dots,n-1\}\bigr{\}}.$
@@
Assume that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4. Then for all $m\in\mathbb{N}_{0}$, $n\in\mathbb{N}$, and an arbitrary choice of $U^{m}_{i}\in\mathcal{U}_{m}$ for each $i\in\{0,1,\dots,n\}$, we have
@@
For each $x\in\bigcap_{i=0}^{n}f_{X}^{-i}(U_{i}^{m}\cap X)=\bigl{\{}z\in U_{0}^{m}\cap X:f_{X}^{i}(z)\in U^{m}_{i}\cap X$ for each $i\in\{1,2,\dots,n\}\bigr{\}}$, there exists $U^{m+n}\in\mathcal{U}_{m+n}$ such that $x\in U^{m+n}$ and $f^{n}(U^{m+n})=U^{m}_{n}$. It follows immediately that $f^{i}(x)\in f^{i}(U^{m+n})\cap U^{m}_{i}\cap X$ and thus $f^{i}(U^{m+n})\cap U^{m}_{i}\cap X\neq\emptyset$ for each $i\in\{1,\,2,\,\cdots,\,n\}$. Therefore $U^{m+n}\in E_{m}(U^{m}_{0},\dots,U^{m}_{n-1};U^{m}_{n})$ and the proof is now complete. ∎
@@
Assume that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4. Then there exists a constant $T_{0}\geqslant 1$ such that for all $n\in\mathbb{N}_{0}$, $k\in\mathbb{N}$, and $W^{n}\in\mathcal{W}_{n}$, there exists $\mathcal{I}\subseteq\mathcal{W}_{n+k}$ with $\operatorname{card}(\mathcal{I})\leqslant(pT_{0})^{k}$ such that
@@
Fix arbitrary integer $n\geqslant N$ and $U^{n}\in\mathcal{U}_{n}$, and consider $\operatorname{B}_{n}^{1}(U^{n})$. We have
@@
for some $U^{0}\in\mathcal{U}_{0}$. Since $U^{n}$ is a preimage under $f^{n-N}$ of $f^{n-N}(U^{n})$, there exists a preimage under $f^{n-N}$ of $U^{0}$, denoted by $U^{n-N}\in\mathcal{U}_{n-N}$, such that $U^{n}\subseteq U^{n-N}$. For each $U^{n+1}\in\mathcal{B}_{n}^{1}(U^{n})\cap\mathcal{U}_{n+1}$, there exists $U^{\prime}\in\mathcal{U}_{n-N}$ which is a preimage of $U^{0}$ such that $U^{n+1}\subseteq U^{\prime}$. Since $U^{n+1}\cap U^{n}\neq\emptyset$ and different preimages of $U^{0}$ are disjoint, we have $U^{\prime}=U^{n-N}$, thus $\operatorname{B}_{n}^{1}(U^{n})\subseteq U^{n-N}$.
Consider $f^{n-N}|_{U^{n-N}}\colon U^{n-N}\rightarrow U^{0}$. By [Deg], this map is at most $p$ to $1$, thus we have
$\operatorname{card}\left(\mathcal{B}_{n}^{1}(U^{n})\right)\leqslant p\cdot\operatorname{card}\bigl{(}\mathcal{B}_{N}^{1}\bigl{(}f^{n-N}(U^{n})\bigr{)}\bigr{)}.$
Let $T_{0}\coloneqq\max\limits_{0\leqslant i\leqslant N}\sup\bigl{\{}\operatorname{card}\bigl{(}\mathcal{B}_{i}^{1}(U)\bigr{)}:U\in\mathcal{U}_{i}\bigr{\}}$, then we have
$\operatorname{card}\left(\mathcal{B}_{n}^{1}(U^{n})\right)\leqslant pT_{0}.$
@@
Thus each element of $\mathcal{W}_{n}$ can be covered by at most $pT_{0}$ many elements in $\mathcal{W}_{n+1}$. Inductively, for every $k\in\mathbb{N}$, each element in $\mathcal{W}_{n}$ can be covered by $(pT_{0})^{k}$ many elements in $\mathcal{W}_{n+k}$. We have established the lemma under the assumption that $n\geqslant N$.
Consider now $n\in\{0,\,\dots,\,N\}$. By the definition of $T_{0}$, each element in $\mathcal{W}_{n}$ can be covered by $T_{0}$ many elements in $\mathcal{W}_{n+1}$. Thus, $T_{0}$ is the constant as desired. ∎
5.3. Proofs of Theorems A and B
In this subsection, we will establish the asymptotic $h$-expansiveness of the map $f|_{X}\colon X\rightarrow X$. Our strategy is to prove the result under slightly stronger assumptions (Theorem 5.14), then we show that for some $N\in\mathbb{N}$, $(f|_{X})^{N}$ satisfies the stronger assumptions, resulting in $h^{*}\left((f|_{X})^{N}\right)=Nh^{*}(f|_{X})=0$.
Here is the main idea of the proof of Theorem 5.14. Recall that we construct in Lemma 5.11 a cover $E_{m}(U_{0}^{m},\cdots,U_{n-1}^{m};U_{n}^{m})$ for each set of the form $\bigcap_{i=0}^{n}f^{-i}{(U_{i}^{m}\cap X)}$ by level-$(m+n)$ good open sets, where $m\in\mathbb{N}_{0}$, $n\in\mathbb{N}$, and each $U_{i}^{m}\in\mathcal{U}_{m}$ is a level-$m$ good open set. Then, we give an upper bound for $\operatorname{card}(E_{m}(U_{0}^{m},\cdots,U_{n-1}^{m};U_{n}^{m}))$ that depends only on m,nm,\,n. Combining Lemma 5.12, we have that for an arbitrary $A\in\bigvee_{i=0}^{n-1}f^{-i}(\mathcal{W}_{m})$, the number of sets in the open cover $\bigvee_{j=0}^{n-1}f^{-j}(\mathcal{W}_{l})$ needed to cover $A$ can be bounded from above, and such an upper bound leads to the conclusion that the topological tail entropy $h^{*}(f)=0$. Bouquets are crucially used in the proof.
First, we note that $X$ is indeed a compact metric space. By $\bigcap_{n\in\mathbb{N}_{0}}\overline{\mathfrak{X}_{n}}, XX$ is closed in $\overline{\mathfrak{X}_{1}}$. Since $\overline{\mathfrak{X}_{1}}$ is compact, $X$ is compact in $\overline{\mathfrak{X}_{1}}$, thus compact with respect to its own subspace topology.
@@ 5.3. Proofs of Theorems A and B
Let $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4. Assume that for each $n\in\mathbb{N}_{0}$, $\mathcal{U}_{n+1}$ is a refinement of $\mathcal{U}_{n}$. Then $f|_{X}\colon X\rightarrow X$ is asymptotically h-expansive.
@@
By Lemma 5.13, $\{\mathcal{W}_{m}\}_{m\in\mathbb{N}_{0}}$ (defined in (4.1)) is a refining sequence of open covers of $X$.
@@
Then we give an upper bound of the cardinality of the set $E_{m}(U^{m}_{0},\dots,U^{m}_{n-1};U^{m}_{n})$. For each $k\in\{0,\,1,\,\dots,\,n\}$ let
(5.14) $\mathcal{L}_{k}\coloneqq\bigl{\{}f^{k}(U):U\in E_{m}(U^{m}_{0},\dots,U^{m}_{n-1};U^{m}_{n})\bigr{\}}.$
@@
We can check that $\{M_{m}\}_{m\in\mathbb{N}_{0}}$ is increasing to $\infty$ and that
@@
$\operatorname{diam}\operatorname{B}_{m}^{n-k}(U^{m}_{k})<C^{\prime\prime}\lambda^{M_{n}}/(2K)\leqslant\delta_{M_{m}},$
where $\delta_{M_{m}}$ is the Lebesgue number of $\mathcal{U}_{M_{m}}$, which is from Proposition 3.10 (iv). Thus, there exists $\widehat{U}_{k}\in\mathcal{U}_{M_{m}}$ such that
@@
For each $0\leqslant k\leqslant n$ and each $0\leqslant t\leqslant\min\{n-k,M_{m}\}$, consider $f^{t}|_{\widehat{U}_{k}}\colon\widehat{U}_{k}\rightarrow f^{t}(\widehat{U}_{k})$. By [Deg], the map is at most $p$ to 1, thus $\operatorname{card}(\mathcal{L}_{k})\leqslant p\cdot\operatorname{card}(\mathcal{L}_{k+t})$. Indeed, if we assume, on the contrary, that $\operatorname{card}(\mathcal{L}_{k})>p\cdot\operatorname{card}(\mathcal{L}_{k+t})$, then there exists at least one collection $\bigl{\{}\widetilde{U}_{1},\,\dots,\widetilde{U}_{p+1}\bigr{\}}\subseteq\mathcal{L}_{k}$ such that $f^{t}\bigl{(}\widetilde{U}_{1}\bigr{)}=\cdots=f^{t}\bigl{(}\widetilde{U}_{p+1}\bigr{)}=\widetilde{U}\in\mathcal{L}_{k+t}$. Since each $\widetilde{U}_{i}$, $1\leqslant i\leqslant p+1$, is a preimage of $\widetilde{U}$ (under $f^{-t}$), $\widetilde{U}_{1},\,\dots,\widetilde{U}_{p+1}$ are pairwise disjoint. Hence, given a point $y\in\widetilde{U}$, by choosing a preimage $x_{i}\in\widetilde{U}_{i}$ for each $0\leqslant i\leqslant p+1$, we obtain a collection of $p+1$ preimages of $y$ in $\widehat{U}_{k}$, which yields a contradiction.
Suppose that $n=s\cdot M_{m}+t$ for some $s,\,t\in\mathbb{N}_{0}$ with $0\leqslant t<M_{m}$. Then, we have
$\operatorname{card}(\mathcal{L}_{0})\leqslant p\cdot\operatorname{card}(\mathcal{L}_{M_{m}})\leqslant\cdots\leqslant p^{s}\cdot\operatorname{card}(\mathcal{L}_{s\cdot M_{m}})\leqslant p^{s+1}\cdot\operatorname{card}(\mathcal{L}_{n}),$
which implies
(5.17) $\operatorname{card}(E_{m}(U^{m}_{0},\dots,U^{m}_{n-1};U^{m}_{n}))\leqslant p^{\frac{n}{M_{m}}+1}.$
Define $\mathcal{I}_{A}\coloneqq\{U\cap X:U\in E_{m}(U^{m}_{0},\dots,U^{m}_{n-1};U^{m}_{n})\}\subseteq\mathcal{W}_{m+n}$. Then $\operatorname{card}(\mathcal{I}_{A})\leqslant p^{\frac{n}{M_{m}}+1}$, and $A\subseteq\bigcup\mathcal{I}_{A}$.
@@
By Lemma 5.12, for each $W^{m+n}\in\mathcal{I}_{A}$ there exists $\mathcal{I}_{W^{m+n}}\subseteq\mathcal{W}_{l+n}$ with $\operatorname{card}(\mathcal{I}_{W^{m+n}})\leqslant(pT_{0})^{l-m}$ such that $W^{m+n}\subseteq\bigcup\mathcal{I}_{W^{m+n}}$, where $p$ is the constant in [Deg] and $T_{0}$ is from Lemma 5.12.
@@
For each $W^{l+n}\in\mathcal{W}_{l+n}$, suppose that $W^{l+n}=U^{l+n}\cap X$, $U^{l+n}\in\mathcal{U}_{l+n}$. By the hypothesis in the statement of Theorem 5.14 that $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ is refining, for each $j\in\{0,1,\dots,n\}$, there exists $U^{l+j}\in\mathcal{U}_{l+j}$ such that $U^{l+n}\subseteq U^{l+j}$. Since $U^{l+j}\subseteq f^{-j}\bigl{(}U^{l}_{j}\bigr{)}$ for some $U^{l}_{j}\in\mathcal{U}_{l}$, we have $U^{l+n}\subseteq U^{l+j}\subseteq f^{-j}\bigl{(}U^{l}_{j}\bigr{)}$. As a consequence
@@
$=$ $\bigcap_{j=0}^{n}\bigl{(}X\cap f^{-j}\bigl{(}U^{l}_{j}\bigr{)}\bigr{)}=\bigcap_{j=0}^{n}f^{-j}\bigl{(}X\cap U^{l}_{j}\bigr{)}\in\bigvee_{j=0}^{n}f^{-j}(\mathcal{W}_{l}).$
By Steps 1 through 3, there exists $\mathcal{J}_{A}\subseteq\bigvee_{j=0}^{n}f_{X}^{-j}(\mathcal{W}_{l})$ with $\operatorname{card}(\mathcal{J}_{A})\leqslant p^{\frac{n}{M_{m}}+1}(pT_{0})^{l-m}$, such that $A\subseteq\bigcup\mathcal{J}_{A}$. It follows that
$\min\biggl{\{}\operatorname{card}\xi_{A}:\xi_{A}\subseteq\bigvee_{j=0}^{n}f_{X}^{-j}(\mathcal{W}_{l}),\ A\subseteq\bigcup\xi_{A}\biggr{\}}\leqslant p^{\frac{n}{M_{m}}+1}(pT_{0})^{l-m}.$
@@
$\qquad\leqslant\log\left(p^{\frac{n}{M_{m}}+1}(pT_{0})^{l-m}\right).$
@@
$\leqslant\lim_{m\to+\infty}\lim_{l\to+\infty}\lim_{n\to+\infty}\frac{1}{n}\log\left(p^{\frac{n-1}{M_{m}}+1}(pT_{0})^{l-m}\right)=\lim_{m\to+\infty}\frac{\log p}{M_{m}}=0.$